OpenStudy (anonymous):

Group 1 y = - 3 halves x + 4 –3x + 2y = 8 Type the first equation from the group you selected and identify the slope and the y-intercept of the equation. Describe how you would graph this line using the slope-intercept method. Be sure to write in complete sentences. Type the second equation from the group you selected and change the equation into slope-intercept form. Be sure to show all of your work. Identify the slope and y-intercept of the equation. Describe how you would graph this line using the slope-intercept method. Be sure to write in complete sentences. In a paragraph of at

OpenStudy (anonymous):

In a paragraph of at least three sentences, explain how the graphs of the two equations in your group are the same and how they are different. Be sure to use complete sentences.

OpenStudy (anonymous):

Well for the first problem, the first equation is in the form y=mx+b where: m=slope b=y intercept So can you pick those out of y=-3/2x+4?

OpenStudy (anonymous):

pick ?

OpenStudy (anonymous):

y=-3/2x+4 is in what is called slope intercept form. y=mx+b m=slope b=y intercept so the slope would be -3/2 and the y intercept would be 4, right?

OpenStudy (anonymous):

okay, sounds right.

OpenStudy (anonymous):

how would I graph this line using the slope-intercept method ?

OpenStudy (anonymous):

Ok. Now, graphing using the slope intercept method when you plot the intercept on your graph, then use the slope to find the next point. So in this case you would plot (0,4) and then go down 3, right 2 to get the next point. plot that one and then draw a line through the two points and you have your line. Understand that?

OpenStudy (anonymous):

barely. im not good at this.

OpenStudy (anonymous):

Slope is "rise/run" -3/2 is down 3 because 3 is negative, and then right 2. 

OpenStudy (anonymous):

I went down 3 units then right 2 units from the y intercept to get another point.

OpenStudy (anonymous):

And then drew a straight line through the two points

OpenStudy (anonymous):

okay i see.

OpenStudy (anonymous):

Ok for the second equation, you can get it in the same form that the first one is in by isolating y on the left side of the equation, then youll be able to pick out the slope and y-intercept just like with the first one

OpenStudy (anonymous):

-3x+2y=8 add 3x to both sides 2y=3x+8 divide both sides by 2 y=3/2x+4 Now that you have that, whats the slope and y intercept?

OpenStudy (anonymous):

3 and 4 i think

OpenStudy (anonymous):

slope=3/2 y intercept=4

OpenStudy (anonymous):

Whatever number thats infront of the x is the slope

OpenStudy (anonymous):

Then you would graph like i did the first one, plot the y intercept and then use the slope to get another point